A DVD is initially at rest so that the
line PQ on the disc’s surface is along
the +x-axis. The disc begins to turn
with a constant z = 5.0 rad/s2.
At t = 0.40 s, what is the angle
between the line PQ and the +x-axis?
1
Expert's answer
2020-08-02T15:17:07-0400

As per the given question,

Initial angular velocity of the disc"(\omega)=0 rad/sec"

Angular acceleration of the disc "(\alpha)=5 rad/sec^2"

angle between the line PQ and x axis="\theta"

time t=0.4sec

"\theta= \omega t+\frac{\alpha t^2}{2}"

Now, substituting the values,

"\theta = 0+\frac{5\times 0.4^2}{2}"

"\Rightarrow \theta =0.4 radian."